OpenStudy (anonymous):

Please help! <3 I learned about Graham's Law today, and on the work sheet I lost it. For the first one it says "A helium atom travels an average of 1000. m/s at 250* C. How fast would an atom of radon travel at the same temperature." Now I know that temp is a measure of Kinetic Energy, and that you always convert the temperature to Kelvin by adding 273. But I am lost as far as where what goes.. please help me!

OpenStudy (anonymous):

@matt101

OpenStudy (matt101):

You actually don't need to worry about temperature in this case. Graham's Law gives the relationship between 2 gases and their speeds (technically "rate of effusion", the rate at which the gases would pass through a small hole). The law itself is this: \[\frac{ R_{1}}{ R_{2}}=\sqrt{\frac{ M_{2} }{ M_{1} }}\] where the R's are the rates of effusion (the speed) of the gases and the M's are the molar masses of the gases. Also be careful because the 1 and 2 are flipped when you go from one side of the equation to the other. The question tells you the speed of the helium gas, and it's easy to figure out its molar mass. You can also figure out the molar mass of radon. You have 3 of the 4 variables, so you just need to plug those into the equation and solve for the unknown - in this case, the speed of the radon atom!

OpenStudy (anonymous):

Youre a live saver!!